A lump of gold is suspected to contain some quantity of aluminum.if the gold has a mass of 450g and is to have a relative density of 5.2, what is the mass of aluminium if the relative density of gold and aluminium are 19.3 and 2.6 respectively.
1
Expert's answer
2019-11-07T09:21:06-0500

Let "x" gm of aluminium is present in it.

Then,

Total volume = volume of aluminium + volume of gold



"\frac{450}{5.2}=\frac{x}{2.6}+\frac{450-x}{19.3}"

"63.22=x(\frac{1}{2.6}-\frac{1}{19.3})"

"x=189.66\ gm"